OpenStudy (anonymous):

Kenneth deposited $750.00 into an account and fifteen years later, the account had earned $787.50 in interest. What simple interest rate did the account earn? Kenneth deposited $750.00 into an account and fifteen years later, the account had earned $787.50 in interest. What simple interest rate did the account earn? Kenneth deposited $750.00 into an account and fifteen years later, the account had earned $787.50 in interest. What simple interest rate did the account earn? Kenneth deposited $750.00 into an account and fifteen years later, the account had earned

OpenStudy (texaschic101):

I = PRT (Interest = principle * rate * time I = 787.50 P = 750 T = 15 now we sub 787.50 = 750 * R * 15 787.50 = 11250 * R -- divide both sides by 11250 787.50/11250 = R .07 = R but this needs to be changed to a percent, so multiply it by 100 .07 x 100 = R 70% = R check.. 787.50 = 750 * .07 * 15 787.50 = 787.50 The account earned 70% simple interest rate